# Cryogenic Thermal Modeling of Microwave High Density Signaling

### v1

@@ III.3 Experimental Measurement of Thermal Conductivity
The thermal conductivity of the outer and inner conductors were each measured using the Thermal Transport Option (TTO) in a DynaCool Physical Property Measurement System (PPMS) by Quantum Design, Inc. Maldonado (1992); zot (a). Using this method, the thermal conductance of the sample is measured and then converted to thermal conductivity based on the temperature sensor lead separation, cross-sectional area, and emissivity.
@@
The TTO measurement setup, shown in Fig. 2, involves the connection of an electrical heater to the free end of a sample and the thermal anchoring of the other end to a variable temperature stage. The heater is operated in a pulsed mode, and two thermometers placed along the sample measure the local temperature. Fitting the rise and fall of the temperature change versus time allows the thermal conductance to be extracted Maldonado (1992). The sample emissivity is a parameter of the thermal model, which includes an estimate of the radiative heat loss.
@@
Samples were measured in the “continuous” measurement mode over a 2 - 300 K temperature range. Inner conductor lead separation was measured to be 2.69 mm, while outer conductor lead separation was 6.96 mm. The inner conductor cross-section was estimated to be 0.0324 mm2, given the reported 0.203 mm outer diameter of the circular inner conductor. The outer conductor cross-section was estimated to be 0.2389 mm2, given the annular shape of the outer conductor cross-section and given the reported 0.86 mm outer diameter of the outer conductor and the reported 0.66 mm outer diameter of the dielectric layer. The cross-sectional dimensions of the cable are summarized in Table 2.
@@
The correction for radiative losses is largest at higher temperatures, even though, in this case, it remains a small correction Krinner et al. (2019). For the emissivity input for these corrections, we selected fixed emissivity values measured at 300 K from the literature, the highest temperature at which we measured thermal conductance. Although emissivity is known to vary with temperature, there is only a slight variation over the range 300 - 50 K where the correction for radiation losses is typically significant compared to the sample thermal conductance. The inner conductor has a 3 $\mu m$ silver coating, and the emissivity is taken to be ϵ=italic-$\epsilon=$ 0.00775, from the measured emissivity at 300 K of a silver coating ($>$ 5 $\mu m$) on a polished SS304L stainless steel surface reported by Woods et al. 2014 Woods et al. (2014). The outer conductor is a cupronickel alloy of 62.5:37.5 Cu to Ni ratio. The emissivity value is taken to be ϵ=italic-$\epsilon=$ 0.0813 based on a curve fit of emissivity versus composition of cupronickel alloy reported in Gelin et al. 2004, where they refer to this emissivity as “thermal emittance” Gelin and Wäckelgård (2004).
@@
Experimental thermal conductivity data for the inner and outer conductors is shown in Fig. 3, along with a PTFE thermal conductivity curve supplied by the NIST Cryogenics Index of Material Properties database Marquardt et al. (2002); Bradley, P.E. et al. (2006).
@@
Figure 3: Thermal conductivity versus temperature for silver-coated cupronickel inner conductor, cupronickel outer conductor, and PTFE. Due to the challenge in accurately modeling the radiative correction above 300 K, the raw data was processed to exclude points above 300 K; the curves shown fit the processed data. The PTFE function displayed is the Teflon model from the NIST Cryogenics Index of Material Properties database Marquardt et al. (2002); Bradley, P.E. et al. (2006).
III.4 Fitting the Thermal Conductivity of SC-086/50-SCN-CN Components
@@
Table 4: Fitting coefficients for thermal conductivity as a function of temperature. “Inner” coefficients fit our $k(T)$ data for the silver-coated C7150 inner conductor of an SC-086/50-SCN-CN cable, and “outer” coefficients fit our $k(T)$ data for the C7150 outer conductor for the same cable. “Inner” and “outer” coefficients were rounded to the nearest thousandth. Data points collected for the inner and outer conductors correspond to temperatures that range from approximately 2.0-297.6 K for the outer conductor and approximately 2.3-292.6 K for the inner conductor. Polynomial fits were used to calculate thermal conductivity for temperatures in the reported ranges for each material and up to 300 K. A linear extension to the origin from the data point corresponding to the lowest temperature in the set of used points was used for modeling thermal conductivities at low temperatures. PTFE coefficients are sourced from the NIST Cryogenics Index of Material Properties database  Marquardt et al. (2002); Bradley, P.E. et al. (2006).
@@
Since our work requires thermal conductivities down to 20 mK, we need to extend the fits below the lower end of the experimental data and the validity of the power series function. Here we assume that the thermal conductivity of materials goes to zero at zero temperature, and we extend the thermal conductivity curve using a straight line that connects the lowest-temperature experimental $k$ data point down to the origin. This procedure likely overestimates the thermal conductivity in this temperature range as the slope of the experimental data near the lowest temperatures is larger than the slope of the straight-line approximation.
@@
Table  4 shows the fitting coefficients $a$ through $i$ for the outer and inner conductors as determined by the least-squares polynomial fit. The coefficients were used to graph the solid black curves of best fit for the inner and outer conductor data in Fig. 3. The NIST PTFE curve was plotted using the coefficients available from the NIST Cryogenics Index of Material Properties database Marquardt et al. (2002); Bradley, P.E. et al. (2006).
III.5 Method of Calculating Static Heat Load
@@
We model the coaxial cable as three parallel thermal conductances, one for each cable component: the outer conductor, dielectric, and inner conductor. The total thermal conductance of the cable is then the sum of each component’s conductance.
@@
While our model does not assume that the temperature profile along any one cable component exactly matches another, it ignores any radial heat transfer across the cable. We justify this simplification by pointing out that the cable’s construction is such that the two conductors are separated by a low thermal conductivity dielectric, an arrangement that limits the radial heat flow. Radial heat flow becomes more relevant in the case of significant Ohmic heating due to a large current flow in the cable where the inner and outer conductors might have very different temperatures.
@@
If the cable components have a constant cross-sectional area and length, the static heat load associated with a given cable component can be calculated by numerically integrating over the thermal conductance for that material as a function of temperature using a fit to the function such as those discussed in Section III.4:
[…]
@@
Where $q_{outer}$, $q_{dielectric}$, and $q_{inner}$ respectively refer to each static heat load associated with the components of a coaxial cable.
@@
Calculated Static Heat load of a Single
SC-086/50-SCN-CN coaxial cable
Fridge Stage Static Heat Load (W)
50K 7.566e-03
4K 3.157e-04
Still 2.373e-06
CP 5.491e-07
MXC 1.132e-08
Table 5: The calculated static heat load of a single SC-086/50-SCN-CN coax cable in a Bluefors XLD1000-SL system at each fridge stage.
@@
The maximum number of Bluefors HDW SC-086/50-SCN-CN cables currently available in a Bluefors XLD1000-SL system is $N$ = 1008 Bluefors (2021). We calculated the fraction of the available cooling power at each stage by taking the ratio between $q_{stage}$ and the cooling power at each stage as listed in Table 1. The results are plotted as a bar chart in Fig. 4 and show that the largest fractional heat load is surprisingly not at the mixing chamber but at the cold plate where the cables take up 64.8% of the available cooling power. There is a sufficient engineering margin for the mixing chamber to reach the base temperature. Using superconducting cables below the 4K stage in place of SC-086/50-SCN-CN would reduce this load dramatically.
Figure 4: The fraction of the estimated cooling powers at each stage taken up by static heat loads of a Bluefors XLD1000-SL system outfitted with the maximum number of cables ($N=1008$). The different refrigerator temperature stages on the x-axis are, as before, denoted as follows: “50K” for the 50K plate, “4K” for the 4K plate, “Still” for the still stage, “CP” for the cold plate, and ”MXC” for the mixing chamber.
@@ IV.1 Overview of Active Heat Loads
All cables carrying an appreciable current will produce an Ohmic heat load within the cryostat. In addition to the resistance of the cable, the heat load in the attenuators used to thermalize the signal line must also be accounted for. Thermalization of the signal line is required to reduce the blackbody radiation present in the cables to a level corresponding to a thermal photon occupation number typically taken to be $\approx$ 10-3, requiring a total attenuation of at least 60 dB Krinner et al. (2019). The effects of thermally generated electrical noise in cables and components on qubit dynamics are examined in Vaaranta et al. (2022); Simbierowicz et al. (2024). Infrared filtering elements are also included in the coax line to prevent radiation from higher temperature stages from propagating through the dielectric layer.
@@
The active load in a coaxial cable is distributed along the cable, but the heat flows predominantly to the lower-temperature end of the cable, where it adds to the static heat load. Due to its smaller cross-sectional area, the inner conductor of the cable has a significantly higher resistance compared with the outer conductor; in our active load model, we only consider the Ohmic heating in the inner conductor. The return currents flow through the much lower-resistance outer conductor and the fridge mounting hardware, which is electrically connected to the grounds of the cables Krinner et al. (2019). In the next section, we measure the inner conductor’s DC electrical resistivity as a function of temperature, which is then used as an input to the active thermal model.
@@
A superconducting quantum computer typically has other sources of heat within the cryostat beyond the active loads in the cables and attenuators, such as semiconductor amplifiers at the 4K stage and TWPA amplifiers at the mixing chamber with a microwave pump signal that is terminated in a 50 $\Omega$ load. The active load thermal models depend on the system configuration and the current in the signal lines. Due to the limited available cooling power, the fraction of available MXC cooling power taken up by the processor is particularly sensitive to the resistance of the cables and the connections to the package holding the quantum processor at the mixing chamber stage. In addition to the limited cooling power, generating heat near the quantum processor creates potential problems as the processor package could become locally hotter than its surroundings unless the processor has an adequate thermal connection to the mixing chamber, something that the available thermometers might not pick up. The static heat load models do not need to include the processor connections at the mixing chamber as all of those cables are isothermal to the MXC plate, and hence, there is no static load below the MXC.
IV.2 Experimental Measurement of DC Resistivity
We measured the DC electrical resistivity of the inner conductor of an SC-086/50-SCN-CN cable between room temperature and 3.8 K as determined by a Cernox temperature sensor. A 0.254 m (${}^{\>\prime\prime}$) length of SC-086/50-SCN-CN coaxial cable was mounted on the 4K stage of a dilution refrigerator by connecting the two ends to two SMA feedthroughs on a KF-40 flange. This guarantees the efficient thermalization of the cable the same way the coaxial cables are anchored to the flanges in conventional coaxial builds in dilution refrigerators. The cable is looped back on itself so that both ends are attached to the 4K stage, ensuring that the length of the cable is isothermal to the 4K plate. The resistance was then measured when cooling down and warming up the fridge using a four-terminal setup through a micro-D connector compatible with the DC looms inside the dilution refrigerators. The resistance data was then scaled by the known length and cross-sectional area to give the resistivity and fitted by an eighth-order polynomial.
@@ IV.2 Experimental Measurement of DC Resistivity
[…]
@@
The resistivity approached a constant value below approximately 10 K, the temperature below which the resistivity is dominated by impurity scattering in the metal, and we observe that this value remains constant down to the base temperature of the fridge when the cable is moved to the MXC. Below the fitting temperature range of T = 3.8 K, we use a constant resistivity value of $\rho=$ 9.928e-9 $\Omega\text{-}m$. We note that RF resistivity will be dominated by the skin effect in the silver coating and will be different from the measured DC resistivity reported here.
IV.3 Active Load Thermal Model
@@
A substantial portion of active heating within a superconducting quantum computer comes from lines with a significant current for long periods, such as flux bias control lines. Meanwhile, lines carrying low, infrequent, short-duration pulses, e.g., one-qubit gate pulses on the XY control lines and readout pulses on the read-in and read-out lines, contribute little to the thermal loads. Flux tunable qubits require a constant offset to set the detuning of their frequency to neighboring qubits and to compensate for both fabrication spread in the Josephson junction, which leads to an offset in Josephson inductance, and for any stray magnetic flux in the SQUID loop. Tunable couplers require a flux bias current to set the operating point where the coupling between adjacent qubits is zero. For the control lines in the active model, we consider only the tunable coupler and qubit flux bias signal line currents.
@@
The mean value of flux bias currents for qubits and tunable couplers in a system depends on the design, and measured values should be used to model a real system. For this work, both the qubit flux bias and tunable coupler flux bias lines are estimated to carry a DC current of I = 0.4 mA. This current is estimated from the mutual inductance between the flux bias line and the qubit of $M=\partial\Phi/\partial I=0.5\Phi_{0}$/mA, where $\Phi_{0}$ is the superconducting flux quantum, and a required flux amplitude of ±$\pm 0.2\Phi_{0}$ Krinner et al. (2019).
@@
The qubit and tunable coupler flux bias line are configured with attenuation at particular fridge stages to ensure the inner signal line is well thermalized to the cooling system Yeh et al. (2017). Each attenuator will dissipate power in the resistors that must be included, feeding only a fraction of the input current to the control line below the attenuator. For this model, we assume an attenuator configuration of a single 20 dB attenuator at the 4K stage of the dilution fridge Krinner et al. (2019) using a “T-pad” attenuator shown in Fig. 6. Matching the 50 $\Omega$ characteristic impedance of the lines along with the required attenuation allows the three resistor values to be determined, which in this case gives $R_{1}=R_{2}=40.91\>\Omega$ and $R_{3}=10.1\>\Omega$ for 20 dB of attenuation. The microwave calibration of attenuators in a typical signal line at cryogenic temperatures has been reported in Simbierowicz et al. (2022).
@@
Figure 6: Circuit diagram for a T-Pad attenuator. Matching the characteristic impedance of the input and output along with the required attenuation allows the three resistor values to be determined, which in this case gives $R_{1}=R_{2}=40.91\>\Omega$ and $R_{3}=10.1\>\Omega$ for 20 dB of attenuation.
@@
Since the output current is smaller than the input current at an attenuator, it is necessary to start with the current required at the lowest temperature stage and work backward up the cable, calculating the input current required to produce the desired output current at each attenuator. We can then use these current calculations, along with measures of cable resistivity and the known resistors in the attenuator, to calculate the active heat load in each line. The assumptions on line currents and attenuator configuration are summarized in Table 7.
@@
Table 7: A summary of the signal lines showing the attenuation configuration and current in each of the various lines, also shown in Fig. 7. An asterisk in the attenuation column indicates that while that line does have attenuators, they are irrelevant for the active load model as the line carries negligible current.
@@
Other heat sources within the cryostat must be added to the total heat loads. TWPA amplifiers Macklin et al. (2015) used in the readout circuit at the MXC stage require an RF pump signal that terminates at a 50 Ohm load. We assume a typical TWPA pump drive of -60 dBm Bluefors (2024), fed by a 20 dB directional coupler. The TWPA pump RF power at the input to the directional coupler will, therefore, be -40 dBm, and the vast majority of that power will go straight through and be dissipated in the 50 Ohm load. The power dissipation allows the RMS current in the pump line at the input to the directional coupler to be calculated to be I = 0.044 mA. Another heat source is from the semiconductor LNA amplifiers that are used in the readout circuit at the 4K stage. A typical LNA that operates over the 4-8 GHz band commonly used in readout dissipates 7.8 mW of power at the 4K stage LowNoiseFactory (2024). The assumptions on TWPA pump and LNA loads are summarized in Table 8.
IV.4 Method of Calculating Active Heat Load
The active heat load is calculated by numerically integrating the resistance value along a cable multiplied by the square of the current. At any stage with an attenuator, the heat load due to the attenuator is given by the sum of the heat loads in the three resistors making up the T-pad attenuator, taking into account the different currents flowing in each. The total active heat load is assumed to flow to the lower fridge stage it is thermally anchored to, so the load quoted for each stage is the active load from the cable above it and the load from the attenuator if one is installed.

### v3

@@ III.2 Experimental Measurement of Thermal Conductivity
A sample of SC-086/50-SCN-CN coaxial cable was deconstructed into its three constituent components: the C7150 cupronickel outer conductor, the PTFE dielectric, and the silver-coated C7150 cupronickel inner conductor. The thermal conductivity of the outer and inner conductors were each measured using the Thermal Transport Option (TTO) in a DynaCool Physical Property Measurement System (PPMS) by Quantum Design, Inc. maldonado1992; zotero-361. Using this method, the thermal conductance of the sample is measured and then converted to thermal conductivity based on the temperature sensor lead separation, cross-sectional area, and emissivity.
@@
The TTO measurement setup, which is shown in Fig. 2, involves the connection of an electrical heater to the free end of a sample and the thermal anchoring of the other end to a variable temperature stage. The heater is operated in a pulsed mode, and two thermometers placed along the sample measure the local temperature. Fitting the rise and fall of the temperature change versus time allows the thermal conductance to be extracted maldonado1992. The sample emissivity is a parameter of the thermal model, which includes an estimate of the radiative heat loss.
@@
Samples were measured over a 2 - 300 K temperature range. Inner conductor lead separation was measured to be 2.69 mm, while outer conductor lead separation was 6.96 mm. The inner conductor cross-section was estimated to be 0.0324 mm2, given the reported 0.203 mm outer diameter of the circular inner conductor. The outer conductor cross-section was estimated to be 0.2389 mm2, given the annular shape of the outer conductor cross-section and given the reported 0.86 mm outer diameter of the outer conductor and the reported 0.66 mm outer diameter of the dielectric layer. The cross-sectional dimensions of the cable are summarized in Table 2.
@@
The correction for radiative losses is largest at higher temperatures, even though, in this case, it remains a small correction krinner2019. For the emissivity input for these corrections, we selected fixed emissivity values measured at 300 K from the literature, the highest temperature at which we measured thermal conductance. Although emissivity is known to vary with temperature, there is only a slight variation over the range 300 - 50 K, where the correction for radiation losses is typically significant compared to the sample thermal conductance. The inner conductor has a 3 $\mu m$ silver coating, and the emissivity is taken to be $\epsilon=$ 0.00775, from the measured emissivity at 300 K of a silver coating (>> 5 $\mu m$) on a polished SS304L stainless steel surface reported by Woods et al. 2014 woods2014. The outer conductor is a cupronickel alloy of 62.5:37.5 Cu to Ni ratio. The emissivity value is taken to be $\epsilon=$ 0.0813 based on a curve fit of emissivity versus composition of cupronickel alloy reported in Gelin et al. 2004, where they refer to this emissivity as “thermal emittance” gelin2004.
@@
Experimental thermal conductivity data for the inner and outer conductors are shown in Fig. 3, along with a PTFE thermal conductivity curve supplied by the NIST Cryogenics Index of Material Properties database marquardt2002; bradleyp.e.2006.
@@
Figure 3: Thermal conductivity versus temperature for silver-coated cupronickel inner conductor, cupronickel outer conductor, and PTFE. Due to the challenge in accurately modeling the radiative correction above 300 K, the raw data was processed to exclude points above 300 K; the curves shown fit the processed data. The PTFE function displayed is the Teflon model from the NIST Cryogenics Index of Material Properties database marquardt2002; bradleyp.e.2006.
III.3 Fitting the Thermal Conductivity of SC-086/50-SCN-CN Components
@@
Table 4: Fitting coefficients for thermal conductivity as a function of temperature. “Inner” coefficients fit our $k(T)$ data for the silver-coated C7150 inner conductor of an SC-086/50-SCN-CN cable, and “outer” coefficients fit our $k(T)$ data for the C7150 outer conductor for the same cable. “Inner” and “outer” coefficients were rounded to seven significant figures, and the full seven significant figures are required to maintain the accuracy of the fits at the high temperature end. Data points collected for the inner and outer conductors correspond to temperatures that range from approximately 2.0-297.6 K for the outer conductor and approximately 2.3-292.6 K for the inner conductor. Polynomial fits were used to calculate thermal conductivity for temperatures in the reported ranges for each material and up to 300 K. A linear extension to the origin from the data point corresponding to the lowest temperature in the set of used points was used for modeling thermal conductivities at low temperatures. PTFE coefficients are sourced from the NIST Cryogenics Index of Material Properties database  marquardt2002; bradleyp.e.2006.
@@
Since our work requires thermal conductivities down to 20 mK, we need to extend the fits below the lower end of the experimental data and the validity of the power series function. Here we assume that the thermal conductivity of materials goes to zero at zero temperature, and we extend the thermal conductivity curve using a straight line that connects the lowest-temperature experimental $k$ data point down to the origin. This procedure likely overestimates the thermal conductivity in this temperature range, as the slope of the experimental data near the lowest temperatures is larger than the slope of the straight-line approximation.
@@
The fitting coefficients $a$ through $i$, as determined by the least-squares polynomial fit for the outer and inner conductors, are listed in Table  4. The full seven significant figures are required to maintain the accuracy of the fits at the high temperature end. We used these coefficients to graph the solid black curves of best fit for the inner and outer conductor data in Fig. 3. We plotted the NIST PTFE curve using the coefficients available from the NIST Cryogenics Index of Material Properties database marquardt2002; bradleyp.e.2006.
III.4 Method of Calculating Static Heat Load
@@
We model the coaxial cable as three parallel thermal conductances, one for each cable component: the outer conductor, dielectric, and inner conductor. The total thermal conductance of the cable is then the sum of the conductance of each component.
@@
While our model does not assume that the temperature profile along any one cable component exactly matches another, it ignores any radial heat transfer across the cable. We justify this simplification by pointing out that the construction of the cable is such that the two conductors are separated by a low thermal conductivity dielectric, an arrangement that limits the radial heat flow. Radial heat flow becomes more relevant in the case of significant Ohmic heating due to a large current flow in the cable, where the inner and outer conductors might have very different temperatures.
@@
If the cable components have a constant cross-sectional area and length, the static heat load associated with a given cable component can be calculated by numerically integrating over the thermal conductance for that material as a function of temperature using a fit to the function such as those discussed in Section III.3:
$q_{component}=\frac{A}{L}\int_{T_{L}}^{T_{H}}k(T)\,dT$ (2)
@@
where $q_{outer}$, $q_{dielectric}$, and $q_{inner}$ respectively refer to each static heat load associated with the components of a coaxial cable. When thermally modeling a line in a dilution fridge, we calculate the static heat load for each temperature stage as the incoming heat from the higher temperature stage minus the heat load being removed to a lower temperature stage.
@@
Calculated Static Heat load of
SC-086/50-SCN-CN coaxial cables
Single Cable NN = 1008 Cables
Fridge Stage Static Heat Load (W) Static Heat Load (W)
50K 6.983e-03 7.039
4K 3.550e-04 3.579e-01
Still 1.963e-06 1.978e-03
CP 6.877e-07 6.932e-04
MXC 1.448e-08 1.460e-05
Table 5: The calculated static heat load of a single SC-086/50-SCN-CN coax cable, and the maximum loading of $N$ = 1008 cables, in a Bluefors XLD1000-SL system at each fridge stage.
@@
The maximum number of Bluefors HDW SC-086/50-SCN-CN cables currently available in a Bluefors XLD1000-SL system is $N$ = 1008 bluefors2021. We calculated the fraction of the available cooling power at each stage by taking the ratio between $q_{stage}$ and the cooling power at each stage as listed in Table 1. The results are plotted as a bar chart in Fig. 4 and show that the largest fractional heat load is surprisingly not at the MXC but rather at the CP, where the cables take up 69.3% of the available cooling power. We had expected that the MXC stage would have the highest fractional heat load since it has the smallest cooling power by a large margin, but upon further reflection, it becomes clear that the architecture that we consider for this paper results in a higher proportional load on the CP than on the MXC. Specifically, the CP has a shorter length of cable run from the higher stage compared to the stages above it, increasing static load from above. Further, the CP has a relatively large temperature differential of 1.4 K to 0.2 K compared with the CP to MXC.
There is a sufficient engineering margin for the MXC to reach the base temperature. Using superconducting cables below the 4K stage in place of SC-086/50-SCN-CN would reduce this load dramatically.
Figure 4: The fraction of the estimated cooling powers at each stage taken up by static heat loads of a Bluefors XLD1000-SL system outfitted with the maximum number of cables ($N=1008$). The different refrigerator temperature stages on the x-axis are, as before, denoted as follows: “50K” for the 50K plate, “4K” for the 4K plate, “Still” for the still stage, “CP” for the cold plate, and “MXC” for the mixing chamber.
@@ IV Active Heat Loads
All cables carrying an appreciable current produce an Ohmic heat load within the cryostat. In addition to the resistance of the cable, the heat load in the attenuators used to thermalize the signal line must also be accounted for. Thermalization of the signal line is required to reduce the blackbody radiation present in the cables to a level corresponding to a thermal photon occupation number, typically taken to be $\approx$ 10-3, requiring a total attenuation of at least 60 dB krinner2019. The effects of thermally generated electrical noise in cables and components on qubit dynamics are examined in vaaranta2022; simbierowicz2024. Infrared filtering elements are also included in the coax line to prevent radiation from higher temperature stages from propagating through the dielectric layer.
@@
The active load in a coaxial cable is distributed along the cable, but the heat flows predominantly to the lower-temperature end of the cable, where it adds to the static heat load. Due to its smaller cross-sectional area, the inner conductor of the cable has a significantly higher resistance compared with the outer conductor; in our active load model, we only consider the Ohmic heating in the inner conductor. The return currents flow through the much lower-resistance outer conductor and the fridge mounting hardware, which is electrically connected to the grounds of the cables krinner2019. In the next section, we measure the DC electrical resistivity of the inner conductor as a function of temperature, which we later use as an input to the active thermal model.
@@
A superconducting quantum computer typically has other sources of heat within the cryostat beyond the active loads in the cables and attenuators, such as semiconductor amplifiers at the 4K stage and TWPAs at the mixing chamber with a microwave pump signal that is terminated in a 50 $\Omega$ load. The active load thermal models depend on the system configuration and the current in the signal lines. Due to the limited available cooling power, the fraction of available MXC cooling power taken up by the processor is particularly sensitive to the resistance of the cables and the connections to the package holding the quantum processor at the mixing chamber stage. In addition to the limited cooling power, generating heat near the quantum processor creates potential problems as the processor package could become locally hotter than its surroundings unless the processor has an adequate thermal connection to the mixing chamber, something that the available thermometers might not pick up. The static heat load models do not need to include the processor connections at the mixing chamber as all of those cables are isothermal to the MXC plate, and hence, there is no static load below the MXC.
IV.1 Experimental Measurement of DC Resistivity
[…]
@@ IV.1 Experimental Measurement of DC Resistivity
Table 6: DC electrical resistivity fit parameters for the inner conductor. The coefficients were rounded to seven significant figures, and the full seven significant figures are required to maintain the accuracy of the fits at the high temperature end. We take the resistivity at temperatures below 3.8 K to be a constant of $\rho=\text{9.928e-9}\>\Omega\text{-m}$.
@@
The resistivity approached a constant value below approximately 10 K, the temperature below which the resistivity is dominated by impurity scattering in the metal, and we observe that this value remains constant down to the base temperature of the fridge when the cable is moved to the MXC. Below the fitting temperature range of T = 3.8 K, we use a constant resistivity value of $\rho=$ 9.928e-9 $\Omega\text{-}m$. We note that RF resistivity is dominated by the skin effect in the silver coating; thus, actual values differ from the measured DC resistivity reported here.
IV.2 Active Load Thermal Model
@@
A substantial portion of active heating within a superconducting quantum computer comes from lines with a significant current for long periods, such as flux bias control lines. Meanwhile, lines carrying low, infrequent, short-duration pulses, e.g., one-qubit gate pulses on the XY control lines and readout pulses on the read-in and read-out lines, contribute little to the thermal loads. Flux-tunable qubits require a constant offset to set the detuning of their frequency to neighboring qubits and to compensate for both fabrication spread in the Josephson junction, which leads to an offset in Josephson inductance, and any stray magnetic flux in the SQUID loop. Tunable couplers require a flux bias current to set the operating point where the coupling between adjacent qubits is zero. For the control lines in the active model, we consider only the tunable coupler and qubit flux bias signal line currents.
@@
The mean value of flux bias currents for qubits and tunable couplers in a system depends on the design, and measured values should be used to model a real system. For this work, we estimate that both the qubit flux bias and tunable coupler flux bias lines carry a DC current of I = 0.4 mA. We arrived at this estimate from the mutual inductance of $M=\partial\Phi/\partial I=0.5\Phi_{0}$/mA between the flux bias line and the qubit, where $\Phi_{0}$ is the superconducting flux quantum, as well as from a required flux amplitude of $\pm 0.2\Phi_{0}$ krinner2019.
@@
The qubit and tunable coupler flux bias line are configured with attenuation at particular fridge stages to ensure the inner signal line is well thermalized to the cooling system yeh2017. Each attenuator dissipates power in the resistors that must be included, feeding only a fraction of the input current to the control line below the attenuator. For this model, we assume an attenuator configuration of a single 20 dB attenuator at the 4K stage of the dilution fridge krinner2019 using a “T-pad” attenuator shown in Fig. 6. Matching the 50 $\Omega$ characteristic impedance of the lines along with the required attenuation allows the three resistor values to be determined, which in this case gives $R_{1}=R_{2}=40.91\>\Omega$ and $R_{3}=10.1\>\Omega$ for 20 dB of attenuation. The microwave calibration of attenuators in a typical signal line at cryogenic temperatures has been reported in simbierowicz2022.
@@
Figure 6: Circuit diagram for a T-Pad attenuator. Matching the characteristic impedance of the input and output, along with the required attenuation, allows the three resistor values to be determined, which in this case gives $R_{1}=R_{2}=40.91\>\Omega$ and $R_{3}=10.1\>\Omega$ for 20 dB of attenuation.
@@
Since the output current is smaller than the input current at an attenuator, it is necessary to start with the current required at the lowest temperature stage and then calculate the input current required at each successive stage to produce the desired output current at each attenuator. We can then use these current calculations, along with measures of cable resistivity and the known resistors in the attenuator, to calculate the active heat load in each line. The assumptions on line currents and attenuator configuration are summarized in Table 7.
@@
Table 7: A summary of the signal lines wherein the attenuation configuration and current in each of the various lines are shown; these can also be found in Fig. 7. An asterisk in the attenuation column indicates that while that line does have attenuators, they are irrelevant for the active load model as the line carries negligible current.
@@
Other heat sources within the cryostat must be added to the total heat loads. TWPAs macklin2015 used in the readout circuit at the MXC stage require an RF pump signal that terminates at a 50 Ohm load. We assume a typical TWPA pump drive of -60 dBm Bluefors-TWPA-2024, fed by a 20 dB directional coupler. The TWPA pump RF power at the input to the directional coupler is therefore -40 dBm, and the vast majority of that power goes straight through and is dissipated in the 50 Ohm load. The power dissipation allows the RMS current in the pump line at the input to the directional coupler to be calculated to be I = 0.044 mA. Another heat source is from the semiconductor Low Noise Amplifiers (LNAs) that are used in the readout circuit at the 4K stage. A typical LNA that operates over the 4-8 GHz band commonly used in readout dissipates 7.8 mW of power at the 4K stage zotero-555. The assumptions on TWPA pump and LNA loads are summarized in Table 8.
IV.3 Method of Calculating Active Heat Load
We calculate the active heat load by numerically integrating the resistance value along a cable multiplied by the square of the current. At any stage with an attenuator, the heat load due to the attenuator is given by the sum of the heat loads in the three resistors making up the T-pad attenuator, taking into account the different currents flowing in each. We assume that the total active heat load flows to the lower fridge stage it is thermally anchored to. Thus, the load quoted for each stage is the active load from the cable above it and the load from the attenuator, if one is installed.